differentiate the following function with respect to x
tan^-1(x – a/x+a)

Grade:12

2 Answers

Rinkoo Gupta
9 years ago

y=tan-1{(x-a)/(x+a)}

dy/dx=1/[1+{(x-a)/(x+a)}2].[{(x+a).1-(x-a).1}/(x+a)2]

=(x+a)2/(x+a)2+(x-a)2.[2a/(x+a)2]

=2a/2(x+a)2

=a/(x+a)2
Thanks & Regards
Rinkoo Gupta
AskIITians faculty
Arun Kumar IIT Delhi
9 years ago
Hi

Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty